Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP:; phone: +1 575-835-7370


News, Travel, Visitors (McMullin)

  • Overall status: In BnA configuration; baselines established; 4-band dipoles in place; correlator good to 58 BlB pairs (some intermittent crate problems on rack 8 under investigation. Stable software system but inching toward next version. New OPT/OST (1.09) deployed; several issues (mostly on the operations side). Many people out this week for sundry travel.
  • Key dates:
    • 07 May - 15 Jun: 4-band science observing
    • 13-31 May: BnA configuration
    • May 18-20, 2011 - Radio Astronomy & the ISM, Durango (JM, VD, CC, MG)
    • May 22-26, 2011 - AAS, Boston
    • May 30, 2011 - NRAO Holiday
  • Staff Travel:
    • LS: 9 May - 20 Jun
    • RP: 16 May - 25 May
    • GM: 25 May - 24 Jun
    • FO: 26 May - 09 Jun
    • FO: 12 Aug - 12 Sep
  • Visitors:
    • Robert Harris (AC982): 24 Apr - 16 Jul
    • Alison Peck (ALMA): 15 May - 17 May
    • Rieks Jager (ALMA): 16 May - 17 May
    • Bob Sault: 16 May - 23 Jun

ALMA News (Peck)

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • Antennas: EA18 out until 27 May.
    • Receivers: this week focus on EA18 (L, S, X Ku).
    • Calibration (Perley): Ongoing investigation into the switched power compression/expansion; testing this week as possible. Bob to support ferreting out the Zeeman testing.
  • WIDAR (Rupen):
    • Notes (Rupen/Sowinski): 3-bit focus; stability tests and long integration this week.
  • Computing (Robnett):

Software status and updates (Butler)

Commissioning Group Reports and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!

Commissioning priorities for the week (McMullin, Chandler)

Single maintenance day week:
  • Mon
    • 0900-1600 Testing
    • 1600-0900(Tues) Science operations and testing
  • Tue:
    • 0900-1700 Testing/software development time
    • 1800-0630: Science operations and testing
  • Wed: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations and testing
  • Thu:
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Fri (not-yet-a-holiday):
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Sat-Sun
    • Science operations + planned testing

Issues Affecting Science Observing:

  • Correlator: [Pokorny/Rupen/McMullin] time mislabeling in SDM data (non-synchronous operation between correlator dumps and CBE can require additional 1/2 integration time padding for online flags).
  • Correlator: [Clark/Carlson] Delay rocking across bands. Implementation pending factor-of-2 reconciliation.
  • Correlator: [Rupen/Carlson w/ Momjian/Sault] V stokes issues; Zeeman observing mode blocked.
  • Array: [Gustaaf] Amplitude wobbles seen on subset of antennas; phase jumps at scan boundaries also seen on subset of antennas (Ka, Q bands only thus far).
  • Array: [All] Zoo of other effects (some data are very good, some still seeing antenna-based IF/Polarization issue).

Testing This week

  • 3-bit samplers (Dhawan+): highest priorities
  • CM/CBE update testing (subarrays+ deconfiguration problems).
  • Spectral Line (Linz/Ott): Spectral dynamic range test (TSPE0010.3923557). Complete.
  • Spectral Line (Goss/Chomiuk): BP stability/repeatability (TSPE0005.3699117). Complete..
  • Correlator: Rupen: S-band shadow testing; This week as possible.
  • Correlator: Durand/McMullin/Pokorny/Butler; review of CRC error events (may involve testing).
  • Correlator: Momjian/Rupen: High frequency Zeeman test
  • Spectral line: Momjian: HI High resolution bandpass testing
  • Array: *; Individual issue sleuthing

Commissioning Groups

    • Adv: Low band (Owen):
      • Notes: Ongoing investigation of the RFI environment; review of last night's data.
  • Calibrator Models
    • Notes: Status report and Ka/S update: 32 and 37 GHz observations available for 3C138, 3C147, 3C48 (3C286 by the end of the week); currently the information is put on the web (rather than added directly to CASA/AIPS; after A configuration data is added, it will be added directly. S band is not yet available.


RSRO B configuration status:

Individual reports (updates)


Demo Science Status


  • Other talks pushed to next week.

Misc Actions

  • [McMullin]: Contact to Paul regarding delay in phased array.
  • [Wrobel-Claire]: Identify L-band continuum projects (first step in determining how to educate these users).
  • [Harland]: OPT 4-band frequency settings; fixed date mode. Fixed in 1.09.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • Development of WIDAR guidelines document (Michael)
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
-- JosephMcMullin - 2011-05-16
Topic revision: r2 - 2011-05-17, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback